# Solving Over-Smoothing in GNNs via Nonlocal Message Passing: Algebraic Smoothing and Depth Scalability

### v1

@@ Abstract
The relationship between Layer Normalization (LN) placement and the over-smoothing phenomenon remains underexplored. We identify a critical dilemma: Pre-LN architectures avoid over-smoothing but suffer from the curse of depth, while Post-LN architectures bypass the curse of depth but experience over-smoothing.
To resolve this, we propose a new method based on Post-LN that induces algebraic smoothing, preventing over-smoothing without the curse of depth. Empirical results across five benchmarks demonstrate that our approach supports deeper networks (up to 256 layers) and improves performance, requiring no additional parameters.
@@
  • Theoretical Characterization: Analysis of LN dynamics and their impact on over-smoothing and the curse of depth.
  • A Principled Solution: A parameter-efficient method that induces algebraic smoothing and avoids over-smoothing and the curse of depth.
@@ 3.2 Analysis of over-smoothing
In this subsection, we empirically demonstrate that the Pre-LN SAN architecture is robust to over-smoothing, whereas the Post-LN counterpart is susceptible to it. We focus on the task of fully supervised node classification. To systematically analyze the impact of depth, we evaluate Pre-LN SANs with varying network depths selected from $\{2,32,64,128,256\}$. Detailed training configurations are provided in the Appendix B.1. Our experiments are conducted on five standard benchmark datasets: Cora, Citeseer, and Pubmed (Yang et al.,, 2016), as well as Coauthor CS and Coauthor Physics (Shchur et al.,, 2019). For all datasets, we employ a random split of $60\%/20\%/20\%$ for training, validation, and testing, respectively.
To characterize the intrinsic dynamical behavior of the architecture prior to optimization, the following table presents the Laplacian energy of the final-layer output for the untrained SAN with Pre-LN.
@@ A.1 Preliminaries
$\frac{\partial\int_{G}|\nabla_{\mu}X_{t}|^{2}d\mu}{\partial t}=$ $\int_{G}-P_{\mu}\Delta_{\mu}X_{t}\cdot proj_{S^{d}}X_{t}d\mu$
$=\sum\limits_{i}(1+\lambda_{i})\lambda_{i}B_{i}C_{i}$

### v2

# Solving Oversmoothing in GNNs via Nonlocal Message Passing: Algebraic Smoothing and Depth Scalability

@@ Abstract
The relationship between Layer Normalization (LN) placement and the oversmoothing phenomenon remains underexplored. We identify a critical dilemma: Pre-LN architectures avoid oversmoothing but suffer from the curse of depth, while Post-LN architectures bypass the curse of depth but experience oversmoothing.
To resolve this, we propose a new method called non-local message passing based on Post-LN that induces algebraic smoothing, preventing oversmoothing without the curse of depth. Empirical results across five benchmarks demonstrate that our approach supports deeper networks (up to 256 layers) and improves performance, requiring no additional parameters.
@@
  • Theoretical Characterization: Analysis of LN dynamics and their impact on oversmoothing and the curse of depth.
  • A Principled Solution: A parameter-efficient method that induces algebraic smoothing and avoids oversmoothing and the curse of depth.
@@ 3.2 Analysis of over-smoothing
In this subsection, we empirically demonstrate that the Pre-LN SAN architecture is robust to over-smoothing, whereas the Post-LN counterpart is susceptible to it. We focus on the task of fully supervised node classification. To systematically analyze the impact of depth, we evaluate Pre-LN SAN with varying network depths selected from $\{2,32,64,128,256\}$. Detailed training configurations are provided in the Appendix B.1. Our experiments are conducted on five standard benchmark datasets: Cora, Citeseer, and Pubmed (Yang et al.,, 2016), as well as Coauthor CS and Coauthor Physics (Shchur et al.,, 2019). For all datasets, we employ a random split of $60\%/20\%/20\%$ for training, validation, and testing, respectively.
To characterize the intrinsic dynamical behavior of the architecture prior to optimization, the following Table 2 presents the Laplacian energy of the final-layer output for the untrained SAN with Pre-LN.
@@ A.1 Preliminaries
$\frac{\partial\int_{G}|\nabla_{\mu}X_{t}|^{2}d\mu}{\partial t}=$ $\int_{G}-2P_{\mu}\Delta_{\mu}X_{t}\cdot proj_{S^{d}}X_{t}d\mu$
$=\sum\limits_{i}2(1-\lambda_{i})\lambda_{i}B_{i}C_{i}$